OpenStudy (anonymous):

A dragster travels 1/4 mile in 6.7 seconds. Assuming acceleration is constant and the dragster is initially at rest, what is its velocity when it crosses the finish line?

OpenStudy (irishboy123):

for constant accceleration from zero, the ave velocity \(\bar v = \dfrac{0 + v_f }{2} = \dfrac{v_f }{2}\) and then \(\bar v = \dfrac{dist }{ time}\)... :p